Determine whether (-8,-5) is a solution of 8x+4y=-8
Answer by checkley77(12844)

(-8,-5) is a solution of 8x+4y=-8
8*-8+4*-5=-5
-64-20=-5
-84=-5 (-8,-5) IS NOT A SOLUTION.